OpenStudy (anonymous):

solve the equation x^2-(6 square root 3)x+24=0, giving your answer in terms of surds, simplified as far as possible.

OpenStudy (tkhunny):

Have you considered the Quadratic Formula? Do you know what a "surd" is?

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

For both questions? Then it is time to use the Quadratic Formula. a = 1 b = -6sqrt(3) c = 24 Let's see what you get.

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

ok i dont get how you plug it in the calculator without it becoming a "math error":/

OpenStudy (tkhunny):

That is what the "surd" part is about. Unless your calculator will do symbolic manipulation, it will not help you with this problem.

OpenStudy (tkhunny):

Discriminant: \[\left(6\cdot\sqrt(3)\right)^{2} - 4\cdot 24\] Can you simplify that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

are the dots addition symbols?

OpenStudy (anonymous):

\[59+12\sqrt{3}\]?

OpenStudy (anonymous):

heeelllpppp

OpenStudy (tkhunny):

Sorry, had to go do my regular job. The dots are multiplication. Please try again. Learning new notation can be frustrating.